OpenStudy (anonymous):

-1[4-(4-95) ] do I began by subtracting inside the parenthesis first? Thank you

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

go from the innermost parens to the outermost.

OpenStudy (anonymous):

is 364?

OpenStudy (anonymous):

no. \[-1[4-(4-95)] = -1[4-(-91)] = -1[4+91] = -1[95] \] \[= -1*95 = -95\]

OpenStudy (anonymous):

4-95 = -91 4-(-91) = 95 -1(95) = -95